# Poisoning Attacks and Defenses in Recommender Systems: A Survey

### v1

@@ 7.1 Limitations
The Benchmarking Shortfall in Secure RS. While the broader field of RS thrives with robust benchmarks and collaborative platforms, secure RS continues to grapple with significant gaps in standardized benchmarking tools. This deficiency hampers the ability to perform consistent, reproducible research across the community, crucial for validating and comparing security enhancements. Although platforms like SDLib, RecAD, and ARLib have been developed to address these needs, they fall short of providing comprehensive coverage across diverse recommendation scenarios and methodologies. These platforms’ limitations complicate the task for researchers attempting to rigorously test and improve secure recommendation algorithms under a variety of complex conditions. The establishment of more inclusive and flexible benchmarking platforms is vital to elevate the experimental rigor and innovation rate in secure RS, ensuring it keeps pace with advancements in general RS.
@@ 7.2 Future Directions
Exploration of Novel Victim Contexts. The emergence of innovative technologies has given rise to a multitude of recommendation scenarios, consequently expanding the landscape of potential threats. Innovations such as multi-modal recommendation Yuan et al. (2023d), LLM-based recommendation Wu et al. (2023b), and self-supervised learning-based recommendation Yu et al. (2023a) are redefining the complexity and capabilities of modern recommender systems. While these technologies offer significant advancements in personalization and accuracy, they also introduce new layers of vulnerability. Research into the security implications of these emerging technologies is still nascent, barely scratching the surface of potential exploit scenarios. Each new recommendation paradigm brings with it unique challenges and security loopholes that differ significantly from those encountered in more traditional RS setups. As we move forward, the focus must be on understanding these novel contexts in depth—identifying their specific vulnerabilities, assessing the risk they pose, and developing robust defensive strategies that can safeguard against both known and unforeseen threats.
@@
Enriching Theoretical Foundation. While various poisoning attack methods have demonstrated the vulnerability of RS, there is a lack of in-depth study on the underlying principles of poisoning attack. For example, a key aspect is the economic balance: the quantity of malicious data injected is limited by economic factors, and similarly, the attack’s effectiveness often has economic goals. This relationship between the attack’s cost and its economic impact is recognized, yet its exact nature and dynamics remain unclear. The poisoning attack domain necessitates robust theoretical foundations addressing such aspects, which would facilitate more streamlined research in poisoning attacks and reduce the dependency on laborious trial-and-error approaches.
Mixture and Long-Term Impact of Vulnerability. Present research in poisoning attacks tends to focus on its immediate impacts, such as the quick injection of poisoning data based on the current data and model configuration. However, this perspective may not fully capture the dynamics of RS. As RS evolve with accumulating data and ongoing model updates, the initial impact of injected poisoning data may diminish or become diluted over time. Therefore, it is essential to investigate not only the immediate effects of a poisoning attack but also to track and analyze the long-term impacts. Understanding how poisoning data interacts with and influences the RS over extended periods can provide deeper insights into the sustained vulnerabilities and resilience of these systems.
Data Neutralization for Defense. An intriguing idea under exploration is the injection of counteractive data into RS to counterbalance the negative impact of poisoning data. This strategy, while proposed in other domains Chan and Ong (2019), has been relatively unexplored in the context of poisoning attacks. The primary challenge in implementing this technique lies in achieving this neutralization effectively and efficiently, without imposing additional burdens on the system or compromising the user experience. The exploration of efficient neutralization techniques promises to be a pivotal step in enhancing the resilience of RS against sophisticated and ever-evolving poisoning strategies.

### v2

@@ 7.1 Limitations
The Benchmarking Shortfall in Secure RS. While the broader field of RS thrives with robust benchmarks and collaborative platforms, secure RS continues to grapple with significant gaps in standardized benchmarking tools. This deficiency hampers the ability to perform consistent, reproducible research across the community, crucial for validating and comparing security enhancements. Although platforms like SDLib [^2] https://github.com/Coder-Yu/SDLib, RecAD [^3] https://github.com/gusye1234/recad, and ARLib [^4] https://github.com/CoderWZW/ARLib have been developed to address these needs, they fall short of providing comprehensive coverage across diverse recommendation scenarios and methodologies. These platforms’ limitations complicate the task for researchers attempting to rigorously test and improve secure recommendation algorithms under a variety of complex conditions. The establishment of more inclusive and flexible benchmarking platforms is vital to elevate the experimental rigor and innovation rate in secure RS, ensuring it keeps pace with advancements in general RS.
@@ 7.2 Future Directions
Exploration of Novel Victim Contexts. The emergence of innovative technologies has given rise to a multitude of recommendation scenarios, consequently expanding the landscape of potential threats. Innovations such as multi-modal recommendation Yuan et al. (2023d), self-supervised learning-based recommendation Yu et al. (2023a), and LLM-based recommendation Wu et al. (2023b) are redefining the complexity and capabilities of modern recommender systems. While these technologies offer significant advancements in personalization and accuracy, they also introduce new layers of vulnerability. Research into the security implications of these emerging technologies is still nascent, barely scratching the surface of potential exploit scenarios. Each new recommendation paradigm brings with it unique challenges and security loopholes that differ significantly from those encountered in more traditional RS setups. As we move forward, the focus must be on understanding these novel contexts in depth—identifying their specific vulnerabilities, assessing the risk they pose, and developing robust defensive strategies that can safeguard against both known and unforeseen threats.
@@
Enriching Theoretical Foundation. While numerous poisoning attack methods have revealed the vulnerabilities of RS, there is a shortage of comprehensive studies on the fundamental principles of poisoning attacks. For instance, a critical aspect is the economic balance: the amount of malicious data injected is constrained by economic factors, and the attack’s effectiveness often aims for economic gains. Although the relationship between the attack’s cost and its economic impact is acknowledged, its precise nature and dynamics remain unclear. The field of poisoning attacks requires robust theoretical foundations to address such aspects, which would enable more efficient research and reduce reliance on labor-intensive trial-and-error methods.
Mixture and Long-Term Impact of Vulnerability. Current research on poisoning attacks primarily focuses on their immediate effects, such as the rapid injection of malicious data based on the existing data and model configuration. However, this approach may not adequately capture the evolving nature of RS. As RS accumulate more data and undergo continuous model updates, the initial impact of injected poisoning data may diminish or become diluted over time. Therefore, it is crucial to investigate not only the short-term effects of a poisoning attack but also to track and analyze its long-term impacts. Understanding how poisoning data interacts with and influences RS over extended periods can provide deeper insights into the sustained vulnerabilities and resilience of these systems.
Data Neutralization for Defense. An intriguing concept under investigation is the injection of counteractive data into recommendation systems RS to mitigate the negative effects of poisoning data. This strategy, while proposed in other domains Chan and Ong (2019), has been relatively unexplored in the context of poisoning attacks. The primary challenge in implementing this technique lies in achieving this neutralization effectively and efficiently, without imposing additional burdens on the system or compromising the user experience. The exploration of efficient neutralization techniques promises to be a pivotal step in enhancing the resilience of RS against sophisticated and ever-evolving poisoning strategies.